OpenStudy (0bucchr):

Laplace transformation of e^(3t+1)

OpenStudy (amistre64):

we can split that up into a constant and a 3t part

OpenStudy (amistre64):

\[\large b^{n+m}=b^n\ b^m\]

OpenStudy (0bucchr):

do continue..

OpenStudy (amistre64):

that was pretty much it; the rest is just using the table of laplace tranforms

OpenStudy (amistre64):

\[L\{e^{3t+1}\}\to\ L\{e^{3t}e\}\to\ e\ L\{e^{et}\}\]

OpenStudy (amistre64):

spose to be a 3t in that last part :/

OpenStudy (0bucchr):

haha thanks. I do know the answer - just testing whether questions for older people have a place here.

OpenStudy (amistre64):

;) if i recall it correctly: L{e^at} = 1/(s-a)

OpenStudy (amistre64):

and if your browser is slowing down, try using google chrome; its spose to be the test browser for the site

OpenStudy (0bucchr):

Thanks. I'm running it now but here in the heart of africa we dont have the fastest connection

OpenStudy (amistre64):

look for the zebra with a radio tag :)

OpenStudy (0bucchr):

haha. Have you heard about Wolfram Alpha?

OpenStudy (0bucchr):

if I write an answer do I get a medal? e/(s-3)